A rational consumer spends all of her income on two goods: Apple and Banana. Suppose the last dollar spent on Apple increased her total utility from 40 utils to 48 utils and the last dollar spent on Banana increased her total utility from 15 utils to 19 utils. If the price of a unit of Apple is 4 Birr, what is the price of a unit of Banana at equilibrium?
1
Expert's answer
2020-10-02T06:47:38-0400

If the last dollar spent on Apple increased her total utility from 40 utils to 48 utils, then MUa = 48 - 40 = 8 utils, and if the last dollar spent on Banana increased her total utility from 15 utils to 19 utils, then MUb = 19 - 15 = 4 utils.

In equilibrium "MUa\/Pa = MUb\/Pb."

If the price of a unit of Apple is 4 Birr, then the price of a unit of Banana at equilibrium is:

"8\/4 = 4\/Pb,"

Pb = 2 Birr.